# Communication-Efficient and Privacy-Adaptable Mechanism for Federated Learning

### v2

@@ III-C Differential Privacy (DP)
Lemma 9 implies that applying a subsampled Gaussian mechanism to a random subsample (obtained by sampling with replacement) of a dataset in a single communication round of the FL framework provides stronger privacy guarantees than applying the Gaussian mechanism to the entire dataset. Since each $k$-th client performs $\tau$ iterations of local SGD (i.e., Equation (3)) on indices sampled uniformly from the dataset $\mathcal{D}^{(k)}$ (i.e., $p_{1}=p_{2}=\ldots=p_{|\mathcal{D}^{(k)}|}=1/|\mathcal{D}^{(k)}|=1/n$, with $n=|\mathcal{D}^{(k)}|$ in (12) or (14), where $p_{j}$ denotes the probability that the data point $\xi_{k,j}$ is chosen at time instance $t$), this procedure is equivalent to subsampling with replacement, which outputs a multiset sample of size $\tau$ following a multinomial distribution $\mathrm{Mult}(\tau;p_{1},\ldots,p_{|\mathcal{D}^{(k)}|})$. [^6] See [21, Lemma 3] for an analogous result for the subsampled Gaussian mechanism under Poisson subsampling. The random multiset sample is then fed into the function $q(\cdot)=(\mathbf{W}_{t+\tau}^{k}-\mathbf{W}_{t})(\cdot)$. Refer to Theorem 12.
@@
[39, Theorem 3] Let $q:\mathcal{X}\rightarrow\mathbb{R}$ be a function with the global $\ell_{1}$-sensitivity $\Delta_{1}:=\sup_{(X,X^{\prime})\in\mathcal{R}}|q(X)-q(X^{\prime})|$ and $\mathcal{F}(X):=q(X)+Z$ with $Z\sim\mathrm{Lap}(0,b)$ be the Laplace mechanism. Then, for any $\epsilon>0$, the privacy profile of $\mathcal{F}$ is given by $\delta_{\mathcal{F}}(\epsilon)=\left[1-e^{\frac{\epsilon-\Delta_{1}/b}{2}}\right]_{+}$, where $\cdot]_{+}:=\max\{\cdot,0\}$.
@@ III-D Problem Setting
Our goal is to construct a mechanism that simultaneously addresses privacy requirements (protecting against honest but inquisitive clients within the FL network or inquisitive data analysts upon public release of the final trained model, see Section III-D1) and compression demands (for lossless uplink channels with limited bandwidth) in a single local model update at each client within the FL framework. Since the distribution of the model parameters and/or the induced gradient is often unknown to the clients, our focus lies in creating a universal mechanism applicable to any random source. Furthermore, we are exploring privacy mechanisms that introduce noise customized to various noise distributions. While Laplace mechanism provides pure DP protection, Gaussian mechanism only provides approximate DP protection with a small failure probability [10, 11]. However, it is well-known that Gaussian mechanism support tractability of the privacy budget in mean estimation [73, 69], an important subroutine in FL. Such schemes can be formulated as mappings from the local update $\mathbf{X}_{t}^{k}\in\mathbb{R}^{m}$ at client $k$ at the time instance $t$ to the estimated update $\hat{\mathbf{X}}_{t}^{k}\in\mathbb{R}^{m}$ at the server, aimed to achieve the following desired properties:
@@ IV CEPAM for FL
In this section, we introduce CEPAM, which utilizes the inherent randomness in probabilistic quantized FL to provide privacy enhancement. (see Algorithm 1). Our approach modifies the schemes in [9] and [21] by replacing universal quantization [57, 58] with RSUQ as outlined in Section III-B. Through the utilization of RSUQ and the layered construction [37, 36], leveraging on the unique quantization error property of RSUQ (refer to Proposition 5), we construct a joint privacy and compression mechanism to address both FL challenges of communication overload and privacy requirement adaptable to any system adjustment simultaneously. This is achieved by applying norm clipping and then LRSUQ to the model updates $\mathbf{X}_{t+\tau}^{k}$ at the end of each FL round, separately for each client. These quantized updates are then transmitted as a set of messages $\{\mathbf{M}_{j}^{k}\}_{j\in\mathcal{N}}$ to the server. The server adds the dithers induced by the shared randomness to the received messages, collects the decoded messages and obtains the estimations $\hat{\mathbf{X}}_{t}^{k}$. Subsequently, the server computes the average of estimated model updates $\frac{1}{K}\sum_{k\in\mathcal{K}}\hat{\mathbf{X}}_{t}^{k}$, which ensures the DP guarantee against the honest but inquisitive clients. The performance of CEPAM will be analyzed in Section V.
@@
3:Initialization: Client $k$ and the server agree on privacy budget $\epsilon>0$ and privacy relaxation $\delta$ for $(\epsilon,\delta)$-DP (or privacy budget $\epsilon>0$ for $\epsilon$-DP), shared seed $s_{k}$, lattice dimension $n$, generator matrix $\mathbf{G}$, privacy-preserving noise $\mathbf{Z}\sim f$ with noise variance $\mathrm{Var}(f)>0$, latent variable $U\sim g(u):=\mu(L_{u}^{+}(f))$, initial model parameter vector $\mathbf{W}_{0}$
@@
9:   Compute $\mathbf{W}_{t+t^{\prime}}^{k}\leftarrow\mathbf{W}_{t{\color[rgb]{0,0,0}+t^{\prime}-1}}^{k}-\eta_{t+t^{\prime}}\nabla F_{k}^{j_{t{\color[rgb]{0,0,0}+t^{\prime}-1}}^{k}}(\mathbf{W}_{t{\color[rgb]{0,0,0}+t^{\prime}-1}}^{k})$
@@
13:  Run subroutine $\mathrm{ENCODE}({\color[rgb]{0,0,0}\tilde{\mathbf{X}}_{t+\tau}^{k}},\gamma,s_{k},(f,g),N)$ $\triangleright$ See Algorithm 2
@@
18:  for $k\in\mathcal{K}$ do
19:   Run subroutine $\mathrm{DECODE}(\{(H_{t+\tau,j}^{k},\mathbf{M}_{t+\tau,j}^{k})\}_{j\in\mathcal{N}},s_{k},g,N)$ $\triangleright$ See Algorithm 3
20:  end for
21:  Compute $\hat{\mathbf{W}}_{t+\tau}\leftarrow\mathbf{W}_{t}+\sum_{k\in\mathcal{K}}p_{k}\hat{\mathbf{X}}_{t+\tau}^{k}$
22:  Set $\mathbf{W}_{t+\tau}\leftarrow\hat{\mathbf{W}}_{t+\tau}$ and broadcast $\mathbf{W}_{t+\tau}$ to all clients, or output $\mathbf{W}_{T}$ if $t+\tau=T$
23:end for
@@
Algorithm 2 $\mathrm{ENCODE}({\color[rgb]{0,0,0}\tilde{\mathbf{X}}_{t+\tau}^{k}},s_{k},(f,g),N)$
@@
5:  Receive $H_{t+\tau,j}^{k}$ and $\mathbf{M}_{t+\tau,j}^{k}$
6:  Use $\tilde{\mathfrak{P}}^{k}$ to sample $\mathbf{V}_{t+\tau,j,1},\ldots,\mathbf{V}_{t+\tau,j,H_{t+\tau,j}^{k}}\stackrel{{\scriptstyle iid}}{{\sim}}\mathrm{Unif}(\mathcal{P})$ $\triangleright$ Initiated by seed $s_{k}$
7:   Compute $\mathbf{Y}_{t+\tau,j}^{k}\leftarrow\beta(U_{t+\tau,j}^{k})(\mathbf{M}_{t+\tau,j}^{k}+\mathbf{V}_{H_{t+\tau,j}^{k}})$ $\triangleright$ Perform decoding, inducing privacy protection
8:end for
9:Collect $\{\mathbf{Y}_{t+\tau,j}^{k}\}_{j\in\mathcal{N}}$ into $\mathbf{Y}_{t+\tau}^{k}$ and set $\hat{\mathbf{X}}_{t+\tau}^{k}\leftarrow\mathbf{Y}_{t+\tau}^{k}$
10:return $\hat{\mathbf{X}}_{t+\tau}^{k}$
Privacy enhancement: The privacy is ensured by LRSUQ executed cooperatively between the clients and the server in compliance with the threat model and the defined privacy guarantee. Through the decoding process, it is ensured that $\{\mathbf{Y}_{t+\tau,j}^{k}\}_{j\in\mathcal{N}}$ comprises noisy estimates of $\{\tilde{\mathbf{X}}_{t+\tau,j}^{k}\}_{j\in\mathcal{N}}$, i.e., the estimated model updates at the server are noisy estimates of the clipped model updates at the clients, thus establishing a privacy mechanism. Refer to Step 7 in Algorithm 3. Specifically, in Section V-A1, we prove that when using the Gaussian mechanism, the global average of estimated model updates satisfies $(\epsilon,\delta)$-DP requirement by specifying appropriate $f$ and $g$, while in Section V-A2, when using the Laplace mechanism, the estimated model updates satisfies $\epsilon$-DP requirement by specifying appropriate $f$ and $g$.
@@ V-A1 Gaussian mechanism
Set $\tilde{U}_{t+\tau,j}^{k}\sim g=\chi_{n+2}^{2}$ and $\tilde{\mathbf{Z}}_{t+\tau,j}^{k}|\{\tilde{U}_{t+\tau,j}^{k}=u\}\sim\mathrm{Unif}(\sigma\sqrt{u}B^{n})$ for every $k,j$, the resulting mechanism, i.e., Step 7 in Algorithm 3, is Gaussian. The average $\frac{1}{K}\sum_{k\in\mathcal{K}}\hat{\mathbf{X}}_{t+\tau}^{k}$ is a noisy estimate of the average of the clipped model updates such that
$\frac{1}{K}\sum_{k\in\mathcal{K}}\hat{\mathbf{X}}_{t+\tau}^{k}=\frac{1}{K}\sum_{k\in\mathcal{K}}{\color[rgb]{0,0,0}\tilde{\mathbf{X}}_{t+\tau}^{k}}+{\color[rgb]{0,0,0}\mathcal{N}\bigg(\mathbf{0},\frac{\sigma^{2}}{K}\mathbf{I}_{m}\bigg)},$ (15)
@@ V-A2 Laplace mechanism
If $g$ follows a Gamma distribution $\mathrm{Gamma}(2,1)$, then $f$ follows a Laplace distribution. [^8] Our CEPAM can be extended to include multivariate Laplace distributions [74] or $t$-distributions [75], which are left for future work. This guarantees that estimated model updates satisfy $\epsilon$-DP. The proof is given in Appendix A-C.
@@
$\hat{\mathbf{X}}_{t+\tau}^{k}={\color[rgb]{0,0,0}\tilde{\mathbf{X}}_{t+\tau}^{k}}+{\color[rgb]{0,0,0}\mathrm{Lap}\left(\mathbf{0},b\mathbf{I}_{m}\right)},$ (17)
@@ V-B Compression
Every client in CEPAM is required to transmit the set of message pairs $(H_{t+\tau,j}^{k},\mathbf{M}_{t+\tau,j}^{k})_{j\in\mathcal{N}}$ per communication round. If we know that $\tilde{\mathbf{X}}_{t+\tau,j}^{k}\in\mathcal{X}$, then we can compress $H_{t+\tau,j}^{k}$ using the optimal prefix-free code [55, 76] for $\mathrm{Geom}(p(u))|\{U_{t+\tau,j}^{k}=u\}$, where $p(u):=\mu(L_{u}^{+}(f))/\mu(\beta(u)\mathcal{P})$, and compress $\mathbf{M}_{t+\tau,j}^{k}|\{U_{t+\tau,j}^{k}=u\}\in\mathcal{M}:=(\mathcal{X}+L_{u}^{+}(f)-\beta(u)\mathcal{P})\cap\mathbf{G}\mathbb{Z}^{n}$ using $H(\lceil\log|\mathcal{M}(U)|\rceil\;|\;U)$ bits. Therefore, the total communication cost per communication round per client is at most $N\cdot\left(H(\mathrm{Geom}(p(U)\;|\;U)+H(\lceil\log|\mathcal{M}(U)|\rceil\;|\;U)\right)$ bits.
@@ VI-A3 Baselines
  • FL+Gaussian: FL with one-dimensional Gaussian mechanism, no compression
  • FL+Gaussian+SDQ: FL with approach that applies one-dimensional Gaussian mechanism followed by scalar SDQ.
  • CEPAM-Gaussian: CEPAM achieves privacy and quantization jointly through LRSUQ to construct Gaussian mechanism, namely CEPAM-Gaussian. In particular, for CEPAM-Gaussian, we evaluate three different cases of LRSUQ that simulate Gaussian noise for dimension $n=1,2,3$, where we use a scaled integer lattice $\alpha\mathbb{Z}^{n}$, where $\alpha=10^{-5}$, [^10] To simplify implementation, we choose to use the integer lattice due to its well-established decoding algorithm [77]. The time complexity of the rejection sampling step can be improved by using a lattice with higher packing density. and the corresponding basic cells $\alpha(-0.5,0.5]$, $\alpha(-0.5,0.5]^{2}$, and $\alpha(-0.5,0.5]^{3}$. We have chosen $\alpha=10^{-5}$ as the precision level for quantization as it aligns with the observed precision level of the model updates during training.

### v3

@@ III-C Differential Privacy (DP)
Lemma 9 implies that applying a subsampled Gaussian mechanism to a random subsample (obtained by sampling with replacement) of a dataset in a single communication round of the FL framework provides stronger privacy guarantees than applying the Gaussian mechanism to the entire dataset. Since each $k$-th client performs $\tau$ iterations of local SGD (i.e., Equation (3)) on indices sampled uniformly from the dataset $\mathcal{D}^{(k)}$ (i.e., $p_{1}=p_{2}=\ldots=p_{|\mathcal{D}^{(k)}|}=1/|\mathcal{D}^{(k)}|=1/n$, with $n=|\mathcal{D}^{(k)}|$ in (12) or (14), where $p_{j}$ denotes the probability that the data point $\xi_{k,j}$ is chosen at time instance $t$), this procedure is equivalent to subsampling with replacement, which outputs a multiset sample of size $\tau$ following a multinomial distribution $\mathrm{Mult}(\tau;p_{1},\ldots,p_{|\mathcal{D}^{(k)}|})$. [^6] See [hasircioglu2023communication, Lemma 3] for an analogous result for the subsampled Gaussian mechanism under Poisson subsampling. The random multiset sample is then fed into the function $q(\cdot)=(\mathbf{W}_{t+\tau}^{k}-\mathbf{W}_{t})(\cdot)$. Refer to Theorem 12.
@@
[Balle2018PrivAmpl, Theorem 3] Let $q:\mathcal{X}\rightarrow\mathbb{R}$ be a function with the global $\ell_{1}$-sensitivity $\Delta_{1}:=\sup_{(X,X^{\prime})\in\mathcal{R}}|q(X)-q(X^{\prime})|$ and $\mathcal{F}(X):=q(X)+Z$ with $Z\sim\mathrm{Lap}(0,b)$ be the Laplace mechanism. Then, for any $\epsilon>0$, the privacy profile of $\mathcal{F}$ is given by $\delta_{\mathcal{F}}(\epsilon)=\left[1-e^{\frac{\epsilon-\Delta_{1}/b}{2}}\right]_{+}$, where $\cdot]_{+}:=\max\{\cdot,0\}$.
@@ III-D Problem Setting
Our goal is to construct a mechanism that simultaneously addresses privacy requirements (protecting against honest but inquisitive clients within the FL network or inquisitive data analysts upon public release of the final trained model, see Section III-D1) and compression demands (for lossless uplink channels with limited bandwidth) in a single local model update at each client within the FL framework. Since the distribution of the model parameters and/or the induced gradient is often unknown to the clients, our focus lies in creating a universal mechanism applicable to any random source. Furthermore, we are exploring privacy mechanisms that introduce noise customized to various noise distributions. While Laplace mechanism provides pure DP protection, Gaussian mechanism only provides approximate DP protection with a small failure probability [Dwork06DP, Dwork14Book]. However, it is well-known that Gaussian mechanism support tractability of the privacy budget in mean estimation [Dong2022GDP, Mironov2017RDP], an important subroutine in FL. Such schemes can be formulated as mappings from the local update $\mathbf{X}_{t}^{k}\in\mathbb{R}^{m}$ at client $k$ at the time instance $t$ to the estimated update $\hat{\mathbf{X}}_{t}^{k}\in\mathbb{R}^{m}$ at the server, aimed to achieve the following desired properties:
@@ IV CEPAM for FL
In this section, we introduce CEPAM, which utilizes the inherent randomness in probabilistic quantized FL to provide privacy enhancement. (see Algorithm 1). Our approach modifies the schemes in [shlezinger2020uveqfed] and [hasircioglu2023communication] by replacing universal quantization [ziv1985universal, zamir1992universal] with RSUQ as outlined in Section III-B. Through the utilization of RSUQ and the layered construction [hegazy2022randomized, wilson2000layered], leveraging on the unique quantization error property of RSUQ (refer to Proposition 5), we construct a joint privacy and compression mechanism to address both FL challenges of communication overload and privacy requirement adaptable to any system adjustment simultaneously. This is achieved by applying norm clipping and then LRSUQ to the model updates $\mathbf{X}_{t+\tau}^{k}$ at the end of each FL round, separately for each client. These quantized updates are then transmitted as a set of messages $\{\mathbf{M}_{j}^{k}\}_{j\in\mathcal{N}}$ to the server. The server adds the dithers induced by the shared randomness to the received messages, collects the decoded messages and obtains the estimations $\hat{\mathbf{X}}_{t}^{k}$. Subsequently, the server computes the average of estimated model updates $\frac{1}{K}\sum_{k\in\mathcal{K}}\hat{\mathbf{X}}_{t}^{k}$, which ensures the DP guarantee against the honest but inquisitive clients. The performance of CEPAM will be analyzed in Section V.
@@
3:Initialization: Client $k$ and the server agree on privacy budget $\epsilon>0$ and privacy relaxation $\delta$ for $(\epsilon,\delta)$-DP (or privacy budget $\epsilon>0$ for $\epsilon$-DP), shared seed $s_{k}$, lattice dimension $n$, generator matrix $\mathbf{G}$, privacy-preserving noise $\mathbf{Z}\sim f$ with noise variance $\mathrm{Var}(f)>0$, latent variable $U\sim g(u):=\mu(L_{u}^{+}(f))$, initial model parameter vector $\mathbf{W}_{0}\in\mathbb{R}^{m}$
@@
9:   Compute $\mathbf{W}_{t+t^{\prime}}^{k}\leftarrow\mathbf{W}_{t+t^{\prime}-1}^{k}-\eta_{t+t^{\prime}-1}\nabla F_{k}^{j_{t+t^{\prime}-1}^{k}}(\mathbf{W}_{t+t^{\prime}-1}^{k})$
@@
13:  Run subroutine $\mathrm{ENCODE}(\tilde{\mathbf{X}}_{t+\tau}^{k},s_{k},(f,g),N)$ $\triangleright$ See Algorithm 2
@@
18:  Receive $\{H_{t+\tau,j}^{k},\mathbf{M}_{t+\tau,j}^{k}\}_{j\in\mathcal{N}}$ from clients
19:  for $k\in\mathcal{K}$ do
20:   Run subroutine $\mathrm{DECODE}(\{(H_{t+\tau,j}^{k},\mathbf{M}_{t+\tau,j}^{k})\}_{j\in\mathcal{N}},s_{k},g,N)$ $\triangleright$ See Algorithm 3
21:  end for
22:  Compute $\hat{\mathbf{W}}_{t+\tau}\leftarrow\mathbf{W}_{t}+\sum_{k\in\mathcal{K}}p_{k}\hat{\mathbf{X}}_{t+\tau}^{k}$
23:  Set $\mathbf{W}_{t+\tau}\leftarrow\hat{\mathbf{W}}_{t+\tau}$ and broadcast $\mathbf{W}_{t+\tau}$ to all clients, or output $\mathbf{W}_{T}$ if $t+\tau=T$
24:end for
@@
Algorithm 2 $\mathrm{ENCODE}(\tilde{\mathbf{X}}_{t+\tau}^{k},s_{k},(f,g),N)$
@@
5:  Use $\tilde{\mathfrak{P}}^{k}$ to sample $\mathbf{V}_{t+\tau,j,1},\ldots,\mathbf{V}_{t+\tau,j,H_{t+\tau,j}^{k}}\stackrel{{\scriptstyle iid}}{{\sim}}\mathrm{Unif}(\mathcal{P})$ $\triangleright$ Initiated by seed $s_{k}$
6:   Compute $\mathbf{Y}_{t+\tau,j}^{k}\leftarrow\beta(U_{t+\tau,j}^{k})(\mathbf{M}_{t+\tau,j}^{k}+\mathbf{V}_{H_{t+\tau,j}^{k}})$ $\triangleright$ Perform decoding, inducing privacy protection
7:end for
8:Collect $\{\mathbf{Y}_{t+\tau,j}^{k}\}_{j\in\mathcal{N}}$ into $\mathbf{Y}_{t+\tau}^{k}$ and set $\hat{\mathbf{X}}_{t+\tau}^{k}\leftarrow\mathbf{Y}_{t+\tau}^{k}$
9:return $\hat{\mathbf{X}}_{t+\tau}^{k}$
Privacy enhancement: The privacy is ensured by LRSUQ executed cooperatively between the clients and the server in compliance with the threat model and the defined privacy guarantee. Through the decoding process, it is ensured that $\{\mathbf{Y}_{t+\tau,j}^{k}\}_{j\in\mathcal{N}}$ comprises noisy estimates of $\{\tilde{\mathbf{X}}_{t+\tau,j}^{k}\}_{j\in\mathcal{N}}$, i.e., the estimated model updates at the server are noisy estimates of the clipped model updates at the clients, thus establishing a privacy mechanism. Refer to Step 6 in Algorithm 3. Specifically, in Section V-A1, we prove that when using the Gaussian mechanism, the global average of estimated model updates satisfies $(\epsilon,\delta)$-DP requirement by specifying appropriate $f$ and $g$, while in Section V-A2, when using the Laplace mechanism, the estimated model updates satisfies $\epsilon$-DP requirement by specifying appropriate $f$ and $g$.
@@ V-A1 Gaussian mechanism
Set $\tilde{U}_{t+\tau,j}^{k}\sim g=\chi_{n+2}^{2}$ and $\tilde{\mathbf{Z}}_{t+\tau,j}^{k}|\{\tilde{U}_{t+\tau,j}^{k}=u\}\sim\mathrm{Unif}(\sigma\sqrt{u}B^{n})$ for every $k,j$, the resulting mechanism, i.e., Step 6 in Algorithm 3, is Gaussian. The average $\frac{1}{K}\sum_{k\in\mathcal{K}}\hat{\mathbf{X}}_{t+\tau}^{k}$ is a noisy estimate of the average of the clipped model updates such that
$\frac{1}{K}\sum_{k\in\mathcal{K}}\hat{\mathbf{X}}_{t+\tau}^{k}=\frac{1}{K}\sum_{k\in\mathcal{K}}\tilde{\mathbf{X}}_{t+\tau}^{k}+\mathcal{N}\bigg(\mathbf{0},\frac{\sigma^{2}}{K}\mathbf{I}_{m}\bigg),$ (15)
@@ V-A2 Laplace mechanism
If $g$ follows a Gamma distribution $\mathrm{Gamma}(2,1)$, then $f$ follows a Laplace distribution. [^8] Our CEPAM can be extended to include multivariate Laplace distributions [andres2013geo] or $t$-distributions [Reimherr2019Elliptical], which are left for future work. This guarantees that estimated model updates satisfy $\epsilon$-DP. The proof is given in Appendix A-C.
@@
$\hat{\mathbf{X}}_{t+\tau}^{k}=\tilde{\mathbf{X}}_{t+\tau}^{k}+\mathrm{Lap}\left(\mathbf{0},b\mathbf{I}_{m}\right),$ (17)
@@ V-B Compression
Every client in CEPAM is required to transmit the set of message pairs $(H_{t+\tau,j}^{k},\mathbf{M}_{t+\tau,j}^{k})_{j\in\mathcal{N}}$ per communication round. If we know that $\tilde{\mathbf{X}}_{t+\tau,j}^{k}\in\mathcal{X}$, then we can compress $H_{t+\tau,j}^{k}$ using the optimal prefix-free code [Golomb1966enc, Gallager1975Geom] for $\mathrm{Geom}(p(u))|\{U_{t+\tau,j}^{k}=u\}$, where $p(u):=\mu(L_{u}^{+}(f))/\mu(\beta(u)\mathcal{P})$, and compress $\mathbf{M}_{t+\tau,j}^{k}|\{U_{t+\tau,j}^{k}=u\}\in\mathcal{M}:=(\mathcal{X}+L_{u}^{+}(f)-\beta(u)\mathcal{P})\cap\mathbf{G}\mathbb{Z}^{n}$ using $H(\lceil\log|\mathcal{M}(U)|\rceil\;|\;U)$ bits. Therefore, the total communication cost per communication round per client is at most $N\cdot\left(H(\mathrm{Geom}(p(U)\;|\;U)+H(\lceil\log|\mathcal{M}(U)|\rceil\;|\;U)\right)$ bits.
@@ VI-A3 Baselines
  • FL+{Gaussian, Laplace}: FL with one-dimensional Gaussian or Laplace mechanism, no compression
  • FL+{Gaussian, Laplace}+SDQ: FL with approach that applies one-dimensional Gaussian or Laplace mechanism followed by scalar SDQ.
  • CEPAM-{Gaussian, Laplace}: CEPAM achieves privacy and quantization jointly through LRSUQ to construct Gaussian or Laplace mechanisms, namely CEPAM-Gaussian or CEPAM-Laplace. In particular, for CEPAM-Gaussian, we evaluate three different cases of LRSUQ that simulate Gaussian noise for dimension $n=1,2,3$, where we use a scaled integer lattice $\alpha\mathbb{Z}^{n}$, where $\alpha=10^{-5}$. [^10] To simplify implementation, we choose to use the integer lattice due to its well-established decoding algorithm [conway1982fast]. The time complexity of the rejection sampling step can be improved by using a lattice with higher packing density. The corresponding basic cells $\alpha(-0.5,0.5]$, $\alpha(-0.5,0.5]^{2}$, and $\alpha(-0.5,0.5]^{3}$. For CEPAM-Laplace using LRSUQ to simulate Laplace noise, we use the scaled integer lattice $\alpha\mathbb{Z}$ with the basic cell (0.5,0.5](-0.5,0.5] for the same $\alpha=10^{-5}$.